# MO-GRPO: Mitigating Reward Hacking of Group Relative Policy Optimization on Multi-Objective Problems

### v1

@@ 3 Group Relative Policy Optimization (GRPO)
GRPO is a reinforcement learning algorithm Shao et al. (2024) which is usually used for online and on-policy learning. For a given state, a policy generates multiple outputs and learns to generate outputs with higher relative reward scores compared to the rest of the outputs. $p_{\mathcal{Q}}$ is the distribution over the initial state (prompt) ($q\sim p_{\mathcal{Q}}$). The policy $\pi_{\theta}\left(\cdot\mid q\right)$ outputs the action (sentence) $o_{g}$ based on the initial state $q$ from the action space. Formally, let $R_{i}$ be the $i$-th reward function, which is a mapping from a prompt-output pair to a scalar value, and assume that there are $K$ reward functions. For each prompt $q$, GRPO samples a group of outputs $\mathbf{o}=\left\{o_{1},o_{2},...,o_{G}\right\}$ from the reference policy $\pi_{\theta_{\text{ref}}}$ and then optimizes the policy model by maximizing the following objective:
$\mathcal{J}(\pi_{\theta})$ $=\mathbb{E}\Biggl[\sum_{g=1}^{G}\frac{1}{G}\frac{1}{|o_{g}|}\frac{\pi_{\theta}\bigl(o_{g}\mid q\bigr)}{\pi_{\theta_{\mathrm{ref}}}\bigl(o_{g}\mid q\bigr)}A_{g}$
$-\beta\text{KL}(\pi_{\theta},\pi_{\theta_{\text{ref}}})\Biggr],$ (1)
@@ 4 Multi-Objective GRPO (MO-GRPO)
$\mathcal{J}(\pi_{\theta})$ $=\mathbb{E}\Biggl[\sum_{g=1}^{G}\frac{1}{G}\frac{1}{|o_{g}|}\frac{\pi_{\theta}\bigl(o_{g}\mid q\bigr)}{\pi_{\theta_{\mathrm{ref}}}\bigl(o_{g}\mid q\bigr)}\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}A_{g}^{\mathrm{MO}}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}$
$-\beta\text{KL}(\pi_{\theta},\pi_{\theta_{\text{ref}}})\Biggr],$ (4)
@@ Appendix B Formal Formulation of GRPO
$\min\!\Bigg(\frac{\pi_{\theta}(o_{g}\mid q)}{\pi_{\theta_{\mathrm{ref}}}(o_{g}\mid q)}\,A_{g},\,$ (9)
$\operatorname{clip}\!\Big(1-\epsilon,\,1+\epsilon,\,\frac{\pi_{\theta}(o_{g}\mid q)}{\pi_{\theta_{\mathrm{ref}}}(o_{g}\mid q)}\Big)\,A_{g}\Bigg)$ (10)
$\qquad\qquad-\beta\,\text{KL}\!\left(\pi_{\theta}\,\middle\|\,\pi_{\theta_{\mathrm{ref}}}\right)\Bigg].$ (11)

### v2

@@ 3 Group Relative Policy Optimization (GRPO)
GRPO is a reinforcement learning algorithm Shao et al. (2024) which is usually used for online and on-policy learning. For a given state, a policy generates multiple outputs and learns to generate outputs with higher relative reward scores compared to the rest of the outputs. $p_{\mathcal{Q}}$ is the distribution over the initial state (prompt) ($q\sim p_{\mathcal{Q}}$). The policy $\pi_{\theta}\left(\cdot\mid q\right)$ outputs the action (sentence) $o_{g}$ based on the initial state $q$ from the action space. Formally, let $R_{i}$ be the $i$-th reward function, which is a mapping from a prompt-output pair to a scalar value, and assume that there are $K$ reward functions. For each prompt $q$, GRPO samples a group of outputs $\mathbf{o}=\left\{o_{1},o_{2},...,o_{G}\right\}$ from the old policy $\pi_{\theta_{\text{old}}}$ and then optimizes the policy model by maximizing the following objective:
$\mathcal{J}(\pi_{\theta})$ $=\mathbb{E}\Biggl[\sum_{g=1}^{G}\frac{1}{G}\frac{1}{|o_{g}|}\frac{\pi_{\theta}\bigl(o_{g}\mid q\bigr)}{\pi_{\theta_{\mathrm{old}}}\bigl(o_{g}\mid q\bigr)}A_{g}\Biggr]$
$-\beta\text{KL}(\pi_{\theta},\pi_{\theta_{\text{ref}}})$ (1)
@@ 4 Multi-Objective GRPO (MO-GRPO)
$\mathcal{J}(\pi_{\theta})$ $=\mathbb{E}\Biggl[\sum_{g=1}^{G}\frac{1}{G}\frac{1}{|o_{g}|}\frac{\pi_{\theta}\bigl(o_{g}\mid q\bigr)}{\pi_{\theta_{\mathrm{old}}}\bigl(o_{g}\mid q\bigr)}\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}A_{g}^{\mathrm{MO}}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Biggr]$
$-\beta\text{KL}(\pi_{\theta},\pi_{\theta_{\text{ref}}})$ (4)
@@ Appendix B Formal Formulation of GRPO
$\min\!\Bigg(\frac{\pi_{\theta}(o_{g}\mid q)}{\pi_{\theta_{\mathrm{old}}}(o_{g}\mid q)}\,A_{g},\,$ (9)
$\operatorname{clip}\!\Big(1-\epsilon,\,1+\epsilon,\,\frac{\pi_{\theta}(o_{g}\mid q)}{\pi_{\theta_{\mathrm{old}}}(o_{g}\mid q)}\Big)\,A_{g}\Bigg)\Bigg]$ (10)
$\qquad\qquad-\beta\,\text{KL}\!\left(\pi_{\theta}\,\middle\|\,\pi_{\theta_{\mathrm{ref}}}\right).$ (11)